OpenStudy (anonymous):

How do you solve nx^2-n-2=0? I keep getting n=2,-1. Can someone explain to me why this isn't right?

hero (hero):

Do you mean n^2 - n - 2 = 0

OpenStudy (anonymous):

yeah lol

hero (hero):

n^2 - n - 2 = 0 n^2-2n+1n - 2 = 0 n(n-2)+1(n-2) = 0 (n+1)(n-2) = 0 n = -1, 2

hero (hero):

Your answer appears to be correct.

OpenStudy (anonymous):

see that's what i got too but my teachers answer on her lecture video is n=2,1

OpenStudy (anonymous):

thanks for your reply i just needed someone to get my same answer = ]